Corrientes Company produces a single product in its Buenos Aires plant that currently sells for 5.40p per unit. Fixed costs are expected to amount to 57,000p for the year, and all variable manufacturing and administrative costs are expected to be incurred at a rate of 2.40p per unit. Corrientes has two salespeople who are paid strictly on a commission basis. Their commission is 10 percent of the sales revenue they generate. (Ignore income taxes.) (p denotes the peso, Argentina’s national currency. Many countries use the peso as their national currency. On the day this exercise was written, Argentina’s peso was worth .104 U.S. dollar.)

Required:

  1. Suppose management alters its current plans by spending an additional amount of 4,000p on advertising and increases the selling price to 6.40p per unit. Calculate the profit on 62,000 units. (Do not round intermediate calculations. Enter your answer in pesos.)
Answer #1

Selling price per unit = 6.40p

All variable manufacturing and administrative costs per unit = 2.40p

Commission to salespersons per unit = 10% of selling price per unit = 6.40p x 10% = 0.64p

Therefore,

Contribution margin per unit = 6.40p - 2.40p - 0.64p = 3.36p

Number of units to be sold = 62,000

Therefore,

Total contribution margin = 3.36p x 62,000 units = 208,320p

Fixed costs for the year = 57,000p

Cost of additional advertising = 4,000p

Therefore,

Total fixed costs for the year = 57,000p + 4,000p = 61,000p

Thus,

Profit on 62,000 units = Total contribution margin - Total fixed costs = 208,320p - 61,000p = 147,320p